OpenStudy (enigma123):

Session 48 Problems PDF: How did they know that the height of the surface was z = 1-x-y? I knew that the region R was bounded by y = -x+1, but how do you go from this equation to the z equation?

OpenStudy (baru):



OpenStudy (baru):

the slant face of the tetrahedron is a 'plane' through the points (0,0,1) (0,1,0) and (1,0,0) so it has the equation in the form ax + by +cz = d substitute each of the three points and we get: c=d b=d a=d thus the equation is dx + dy + dz=d or x+y+z=1 => z=1-x-y

OpenStudy (phi):

Another way, if you know cross products. if we call the 3 points a, b, c then form the vectors (for example) b-a, and c-a now form the cross product and you will find the vector normal to the plane once we have the normal we solve for the constant: N dot P = c thus: (0,0,1) - (0,1,0) = (0, -1, 1) (1,0,0) - (0,1,0)= (1, -1, 0) cross product of (0,-1,1) X (1, -1, 0) = (1, 1, 1) now we can do <1,1,1> dot <x,y,z> = c put in any point on the plane, e.g. (1, 0 , 0) we get 1+0+0= c and the equation of the line is x +y + z = 1 and finally z= 1-x-y more work than baru's approach (where we have lots of zeros), but it might be easier if we are given "harder" points.